OpenStudy (anonymous):

(3a4)3 (-2x3y)4 (-mn8)2 (3x3y)2(-4x2y4)3

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

each one is a new equation

OpenStudy (anonymous):

1. take 3 and multiply it by 3a and then by 4 2.take 4 and multiply it by 2x and then by 3y 3.take 2 and multiply it by -mn8 4.take 2 and multiply it by 3x then 3y then take 3 and multiply it by -4x then by 2y then by 4 4

OpenStudy (anonymous):

1=9a12 2=-8x12y 3=-2mn16 4=-6x

OpenStudy (mathstudent55):

\( (a^m)^n = a^{mn}\) 1. \( (3a^4)^3 = (3^1a^4)^3 = 3^{3 \times 1}a^{4 \times 3} = 3^3a^{12} = 27a^{12}\)

OpenStudy (mathstudent55):

2. \((-2x^3y)^4 = (-2)^4(x^3)^4 y^4 = 16x^{12}y^4\)

OpenStudy (anonymous):

so is that the final answer

OpenStudy (mathstudent55):

Those are the final answers for 1 and 2.